Consider a palletizer at a bottling plant that has a first cost of $141,000, operating and maintenance costs of $16,300 per yeconomics analysis
Answer #1

Given,

Study period = 34 years

Interest rate = 8.50% per year

The first cost of the palletizer = $141,000

Annual Operating and maintenance costs = -$16,300

(P/A, i, N) = (1+i)-1 i(i+1)

Present worth of the annual O & M costs = -$16,300 x (P/A, 8.50%, 34) = $16,300 x [(1+0.085)34-1]/[0.085(0.085+1)34] = -$179,793

(P/F, i, N) = (1+i)

Salvage value at the end of 34 years = +$23,500

Present Worth of the salvage value = +$23,500 x (P/F, 8.50%, 34) = $23,500 x 1/(1+0.085)34 = +$1,467

Net Present Worth of the investment = First cost +PW of the annual O&M costs + PW of the Salvage Value

= -141,000 -$179,793 + $1,467 = -$319,326

Ans: $319,326